# The Frequency Response of Networks as Open Systems

### v1

@@ IV.2 Properties of Controllability Gramian
IV.2.1 Trace of Continuous-Time Controllability Gramian
If the largest real part of the eigenvalues of the matrix $A$ is large in magnitude, the trace of the continuous-time Gramian will be proportional to the number of input nodes $m$, i.e.,
@@ IV.2.1 Trace of Continuous-Time Controllability Gramian
[…]
@@
An example of this is shown in Fig. 7, where $\text{Tr}(W_{c})$ is evaluated for selected real networks. We see that when the largest real part of the eigenvalues of the adjacency matrices is large in magnitude (panel a), $\text{Tr}(W_{c})\approx m/(2|a|)$, and when the magnitude is small (panel b), the trace no longer has a linear relationship with $m$.
FIG. 7: Linear relationship between the trace and the number of input nodes. The trace of the infinite-horizon continuous-time controllability Gramian $\text{Tr}(W_{c})$ for different real networks vs their number of input nodes $m$. In panels a and b, the adjacency matrices are scaled such that all networks have the largest real part of the eigenvalues of $-10^{3}$ and $-1$, respectively. The dashed black line demonstrates the linear relationship in Eq. (20): $\text{Tr}(W_{c})=m/2|a|=m/2(10^{3})$. The same dashed black line is maintained in Panel b for easier comparison.
IV.2.2 Modularity
It was shown that properties such as the trace of the output controllability Gramian have the modularity property [64]. Modularity is analogous to linearity, which means that each element of a subset gives an independent contribution to the function value. In particular case of the trace of the output controllability Gramian, it is shown that the contribution of each column of the matrix $B$ to the trace is independent of the others, so one can check the contribution of each column individually and select the columns that provide desired contributions (e.g., the columns that minimize/maximize the trace.)
@@ IV.2.2 Modularity
In [64], it was shown that the actuator selection for control of linear systems can be formulated as a set optimization.
@@
Modular set functions are similar to linear functions and have the property that each element of a subset gives an independent contribution to the function value. Consequently, if $f$ is modular, the optimization problem (23) is easily solved by simply evaluating the set function for each element, sorting the result, and then choosing the top $k$ individual elements from the sorted list to obtain the best subset of size $k$. It was shown that properties such as the trace of the continuous-time controllability Gramian have the modularity property, while the smallest eigenvalue of the controllability Gramian does not [64].
The trace of the output controllability Gramian $CW_{c}C^{\top}$ also has a modularity property with respect to the rows of the output matrix $C$. This means that $\text{Tr}(CW_{c}C^{\top})=\sum_{j=1}^{N}C_{j,:}W_{c}C_{j,:}^{\top}$, where $C_{j,:}\in\mathbb{R}^{1\times n}$ is row $j$ of the matrix $C$. To see this, we use the cyclicity and linearity of the trace and write
@@
One interesting question is how to select a set of $k$ nodes in the network as input nodes so that the trace of the output controllability Gramian is maximized or minimized. Through the modularity property of the trace, the trace is recorded when each node is set as the input node. Then, $k$ nodes that have resulted in the maximum/minimum trace are selected as the optimal choices for input nodes.
IV.2.3 Properties of the trace of the Output Controllability Gramian
@@
The proof follows from these two facts: (i) the output controllability Gramian with output set $\mathcal{O}_{i}$ is equal to a minor of the controllability Gramian obtained by selecting the rows/columns of the controllability Gramian corresponding to the nodes in $\mathcal{O}_{i}$ [9] and (ii) the trace of a square matrix is equal to the sum of the elements on the main diagonal. From (27) and (29) it follows trivially that
@@ Acknowledgements
We acknowledge support from grants AFOSR FA9550-24-1-0214 and Oak Ridge National Laboratory 006321-00001A.

### v2

@@ IV.4 Properties of Controllability Gramian
IV.4.1 Trace of Continuous-Time Controllability Gramian
We proceed under the assumption that all the eigenvalues of the matrix $A$ have negative real part. We see that an important role is played by the largest real part of the eigenvalues $a=\max_{i}\Re(\lambda_{i}(A))<0$. If $a$ is large in magnitude, the trace of the continuous-time Gramian will be proportional to the number of input nodes $m$, i.e.,
@@ IV.4.1 Trace of Continuous-Time Controllability Gramian
$\begin{split}AW_{c}+W_{c}A^{\top}+BB^{\top}=&\,(aI+\Delta)W_{c}+W_{c}(aI+\Delta)^{\top}\\ &\,+BB^{\top}\\ =&\,2aW_{c}+\Delta W_{c}+W_{c}\Delta^{\top}+BB^{\top}.\end{split}$ (26)
@@
An example of this is shown in Fig. 8, where $\text{Tr}(W_{c})$ is evaluated for selected real networks. We see that when $a$, the largest real part of the eigenvalues of the matrix $A$ is large in magnitude (panel a), $\text{Tr}(W_{c})\approx m/(2|a|)$, and when the magnitude is small (panel b), the trace no longer has a linear relationship with $m$.
FIG. 8: Linear relationship between the trace and the number of input nodes. The trace of the infinite-horizon continuous-time Gramian $\text{Tr}(W_{c})$ for different real networks vs their number of input nodes $m$. In panels a and b, the matrices AA are shifted so that for all networks the largest real part eigenvalues is $10^{3}$ and 1-1, respectively. The dashed black line demonstrates the linear relationship in Eq. (25): $\text{Tr}(W_{c})=m/2|a|=m/2(10^{3})$. The same dashed black line is maintained in Panel b for easier comparison.
IV.4.2 Modularity
It was shown that properties such as the trace of the output Gramian have the modularity property [79]. Modularity is analogous to linearity, which means that each element of a subset gives an independent contribution to the function value. In particular case of the trace of the output Gramian, it is shown that the contribution of each column of the matrix $B$ to the trace is independent of the others, so one can check the contribution of each column individually and select the columns that provide desired contributions (e.g., the columns that minimize/maximize the trace.)
@@ IV.4.2 Modularity
In [79], it was shown that the actuator selection for the control of linear systems can be formulated as a set optimization.
@@
Modular set functions are similar to linear functions and have the property that each element of a subset gives an independent contribution to the function value. Consequently, if $f$ is modular, the optimization problem (28) is easily solved by simply evaluating the set function for each element, sorting the result, and then choosing the top $k$ individual elements from the sorted list to obtain the best subset of size $k$. It was shown that properties such as the trace of the continuous-time Gramian have the modularity property, while the smallest eigenvalue of the Gramian does not [79].
The trace of the output Gramian $CW_{c}C^{\top}$ also has a modularity property with respect to the rows of the output matrix $C$. This means that $\text{Tr}(CW_{c}C^{\top})=\sum_{j=1}^{N}C_{j,:}W_{c}C_{j,:}^{\top}$, where $C_{j,:}\in\mathbb{R}^{1\times n}$ is row $j$ of the matrix $C$. To see this, we use the cyclicity and linearity of the trace and write
@@
One interesting question is how to select a set of $k$ nodes in the network as input nodes so that the trace of the output Gramian is maximized or minimized. Through the modularity property of the trace, the trace is recorded when each node is set as the input node. Then, the $k$ nodes that have resulted in the maximum/minimum trace are selected as the optimal choices for input nodes.
IV.4.3 Properties of the trace of the Output Controllability Gramian
@@
The proof follows from these two facts: (i) the output Gramian with output set $\mathcal{O}_{i}$ is equal to a minor of the Gramian obtained by selecting the rows/columns of the Gramian corresponding to the nodes in $\mathcal{O}_{i}$ [9] and (ii) the trace of a square matrix is equal to the sum of the elements on the main diagonal. From (32) and (34) it follows trivially that
@@
IV.5 Dynamical models for each network type
The matrix $A$ considered in this work is the Jacobian of the network dynamics close to a stable fixed point, for each type of network. Therefore, it depends on both the specific structure of the network and the dynamics that it supports. Here, we describe the dynamical models we used for each type of network dynamics considered in this study.
IV.5.1 Food webs
We employ the generalized Lotka-Volterra model, which is a well-known model of competition and mutualism among interacting species [18, 80, 81, 82]. The governing equations of this model are:
$\dfrac{dz_{i}(t)}{dt}=z_{i}(t)\left(r_{i}-s_{i}z_{i}(t){\color[rgb]{0,0,0}+}\sum_{j\neq i}\tilde{A}_{ij}z_{j}(t)\right),$ (36)
$i=1,\ldots,N$, where $z_{i}(t)$ is the population of species $i$, $r_{i}$ are the intrinsic rates of birth if $r_{i}>0$, or death if $r_{i}<0$. The positive constants $s_{i}$ represent the finite carrying capacity of the ecosystem (limited resources) and prevent species $i$ from growing indefinitely. The interaction among species is given by the adjacency matrix $\tilde{A}=[\tilde{A}_{ij}]$, which has zeros on its main diagonal, i.e., $\tilde{A}_{ii}=0,\forall i$.
Following [18], we assume $s_{i}=1,\forall i$, for simplicity. For a proper choice of $r_{i}$, the point $z_{i}^{*}=1,\forall i$, is the fixed-point of the system (if $r_{i}=\sum_{j\neq i}\tilde{A}_{ij},\forall i$). Under these conditions, the Jacobian matrix $A=[A_{ij}]$ evaluated at the fixed-point $x_{i}^{*}=1$ becomes
$A_{ij}=\begin{cases}-\sum_{k\neq i}\tilde{A}_{ik},&\quad\text{if}\ \ i=j,\\ \\ -\tilde{A}_{ij},&\quad\text{if}\ \ i\neq j.\end{cases}$ (37)
The Jacobian $A$ has all non-positive real-part eigenvalues for networks with non-negative weights (this follows from the Gershgorin circle theorem). In practice, it may sometimes happen that the matrix $A$ has zero eigenvalues, which is undesired when calculating the Gramian (the Gramian will not converge). Therefore, we apply the spectral shift $A\leftarrow A-cI$ and properly choose $c$ so that the largest real part eigenvalue of the matrix $A$ is 1-1. The input nodes are identified and the matrix $B$ is constructed as detailed in Supplementary Note 4A. We then perform our analysis for the pair $(A,B)$, where the matrix $B$ is constructed using knowledge of the particular selection of the input nodes from the empirical datasets.
IV.5.2 Connectomes
Following [83, 84], we use the Kuramoto model [36] to describe the interaction dynamics within connectome networks,
$\dfrac{d\theta_{i}(t)}{dt}=\omega_{i}-\frac{K}{N}\sum_{j=1}^{N}\tilde{A}_{ij}\sin(\theta_{i}(t)-\theta_{j}(t)),$ (38)
$i=1,\ldots,N$, where $\theta_{i}\in(-\pi,\pi]$ and $\omega_{i}\in\mathbb{R}$ correspond, respectively, to the phase and the natural frequency at node $i$ . The $N$-dimensional matrix $\tilde{A}=[\tilde{A}_{ij}]$ is the adjacency matrix, describing the network topology, and $N$ is the number of nodes. When $\kappa=\frac{K}{N}$ is large enough, the system typically reaches a phase-locking state where all the phases evolve at the same angular frequency given by $\Omega=N^{-1}\sum_{j=1}^{N}\omega_{j}$ . The phase-locked state is defined as when $d\theta_{i}(t)/dt=\Omega,\forall i$. Therefore, the phased-locked state for oscillator $i$ is $\theta_{i}^{*}(t)=\theta_{i}^{*}(0)+\Omega t$.
The Jacobian of the system $A=[A_{ij}]$ linearized about the phase-locked state is,
$A_{ij}=\begin{cases}-\sum_{k}\tilde{A}_{ik}\cos(\theta_{i}^{*}(0)-\theta_{k}^{*}(0)),&\quad\text{if}\ \ i=j,\\ \\ \tilde{A}_{ij}\cos(\theta_{i}^{*}(0)-\theta_{j}^{*}(0)),&\quad\text{if}\ \ i\neq j.\end{cases}$ (39)
In this case, the Jacobian $A$ has some of the entries in its rows equal zero, thus it has at least one eigenvalue equal to 0. Since our methodology requires asymptotic stability, we apply the spectral shift $A\leftarrow A-cI$ and properly choose $c$ so that the largest real part eigenvalue of the matrix $A$ is 1-1. The input nodes are identified and the matrix $B$ is constructed as detailed in Supplementary Note 4B. We then perform our analysis for the pair $(A,B)$, where the matrix $B$ is constructed from knowledge of the particular selection of the input nodes derived from the empirical datasets.
In our simulations, we randomly select the natural frequencies $\omega_{i}$ from the standard uniform distribution (between 0 and 1). We set $K=N$, so that $\kappa=1$.
IV.5.3 Power grids
We consider a simplified swing dynamics for the voltage angles, including losses on the transmission lines. The dynamics obeys the equations [33],
$\dfrac{d\theta_{i}(t)}{dt}$ $=P_{i}-\sum_{j\neq i}B_{ij}\sin(\theta_{i}-\theta_{j})+G_{ij}[1-\cos(\theta_{i}-\theta_{j})],$ (40)
$i=1,\ldots,N$, where $\theta_{i}\in(-\pi,\pi]$ is the voltage phase and $P_{i}$ is the injected/consumed power at bus (node) $i$ . The transmission line between buses $i$ and $j$ have susceptance $B_{ij}$, and conductance $G_{ij}$ . The Jacobian matrix $A$ is then obtained by linearizing the dynamics around a stable operational state of the grid which corresponds to a phase-locked state similar to Sec. IV.5.2. Each fixed point depends on the topology of the grid and the dispatch of injected/consumed power $P_{i}$ . The entries of the Jacobian are,
$A_{ij}=\begin{cases}-\sum_{k\neq i}\tilde{A}_{ik}\cos(\theta_{i}^{*}-\theta_{k}^{*}+\gamma_{ij}),&\quad\text{if}\ \ i=j,\\ \\ -\tilde{A}_{ik}\cos(\theta_{i}^{*}-\theta_{j}^{*}+\gamma_{ij}),&\quad\text{if}\ \ i\neq j,\end{cases}$ (41)
where $\tilde{A}_{ij}=\sqrt{B_{ij}^{2}+G_{ij}^{2}}$ , and $\gamma_{ij}=\arctan(-G_{ij}/B_{ij})$ . As the rows of the matrix $A$ in (41) sum to zero, it has at least one vanishing eigenvalue. Since our methodology requires asymptotic stability, we apply the spectral shift $A\leftarrow A-cI$ and properly choose $c$ so that the largest real part eigenvalue of the matrix $A$ is 1-1. The input nodes are identified and the matrix $B$ is constructed as detailed in Supplementary Note 4C. We then perform our analysis for the pair $(A,B)$, where the matrix $B$ is constructed from knowledge of the particular selection of the input nodes derived from the empirical datasets.
IV.5.4 Genetic and Pathway networks
We use the regulatory model (Michaelis-Menten model [78]) to capture pathways and genetic dynamics. The model obeys the equations,
$\dfrac{dz_{i}(t)}{dt}=-f_{i}z_{i}^{a}(t)+\kappa\sum_{j=1}^{N}\tilde{A}_{ij}\dfrac{z_{j}^{h}(t)}{1+z_{j}^{h}(t)},$ (42)
$i=1,\ldots,N$, where $z_{i}(t)$ is the gene expression at time $t$, the local dynamics $f_{i}z_{i}^{a}(t)$ captures biochemical processes (where $a$ depends on the process), the matrix $\tilde{A}=[\tilde{A}_{ij}]$ is the weighted adjacency matrix of the network, and the term $z_{j}^{h}(t)/(1+z_{j}^{h}(t))$ describe genetic activation, where $h$ modulates the saturation of the term and is associated with the level of cooperation in gene regulation. For further details of the model and its nonlinear analysis, see [27, Supplementary Section 4.2].
The Jacobian of the system linearized about the active fixed-point (a fixed-point other than the origin) $\boldsymbol{z}^{*}=[z^{*}_{i}]$ is $A=[A_{ij}]$ where
$A_{ij}=\begin{cases}-f_{i}a{z_{i}^{*}}^{a-1},&\quad\text{if}\ \ i=j,\\ \\ \kappa\tilde{A}_{ij}\dfrac{h{z^{*}}_{j}^{h-1}}{(1+{z^{*}}_{j}^{h})^{2}},&\quad\text{if}\ \ i\neq j.\end{cases}$ (43)
As with other models, we apply the spectral shift $A\leftarrow A-cI$ and properly choose $c$ so that the largest real part eigenvalue of the matrix $A$ is 1-1. We then perform our analysis for the pair $(A,B)$, where the matrix $B$ is constructed from knowledge of the particular selection of the input nodes derived from the empirical datasets, see Supplementary Notes 4D and 4E.
In our simulations, we set $a=2$, $h=1$, $\kappa=1$, and sample $f_{i}$ from the standard uniform distribution (between 0 and 1). With this choice of parameters, an active fixed-point (a fixed-point other than the origin) exists and is stable.
@@ Acknowledgements
We acknowledge support from grants AFOSR FA9550-24-1-0214 and Oak Ridge National Laboratory 006321-00001A. Wai Lim Ku was supported by the National Institutes of Health grant K22HL153477